OpenStudy (anonymous):

Write the equation in its equivalent logarithmic form. 23 = x

OpenStudy (mathmate):

I suppose you mean \( 2^3 = x \) Hint: Definition of logarithm to base b is If \(b^x\ =\ c\) then \( x\ =\ log_b\ c \)

OpenStudy (anonymous):

? SO LOST

OpenStudy (anonymous):

its just says x

OpenStudy (mathmate):

Is the question 23=x (as you wrote it) or does it actually read \( 2^3\ =\ x \) i.e. the 3 is smaller and is higher than the 2?

OpenStudy (mathmate):

Compare with the definition of logarithm and deduce the results accordingly.